OpenStudy (anonymous):

solve for the derivative og the folowing function implicitly x^3 - 3x^2y + 3xy^2 = 12 help me to solve this.

OpenStudy (anonymous):

@UnkleRhaukus help please!

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

@wio help

OpenStudy (anonymous):

@ajspeller help

OpenStudy (anonymous):

@butterflydreamer help!

OpenStudy (anonymous):

Why did you close and reopen?

OpenStudy (anonymous):

\[ x^3 - 3x^2y + 3xy^2 = 12\\ d(x^3 - 3x^2y + 3xy^2) = d(12)\\ d(x^3) - 3~d(x^2y) + 3~d(xy^2) = 0 \]

OpenStudy (ajspeller):

 I believe this is correct